OpenStudy (anonymous):

find a real number between 5 and the square root of 36

rebeccaxhawaii (rebeccaxhawaii):

square root of 36 is 6 so 5-6

OpenStudy (faiqraees):

Any number between 5 and.....?

OpenStudy (anonymous):

\[5=\sqrt{5^2}=\sqrt{25}\] between \[\sqrt{25}~and~\sqrt{36}~you can~write~many~real~numbers.\] or you can write infinite real numbers between 5 and 6.

jhonyy9 (jhonyy9):

but like an attention dont forgot that sqrt36 = +/- 6 so in this case how ?